# Extending Douglas–Rachford Splitting for Convex Optimization

### v1

@@ 7 Operator Composition Interpretation
$R_{A}^{\alpha,\beta}=\left(1+\frac{\beta}{\alpha}\right)\frac{1}{2}\left(\operatorname*{Id}+N\right)-\frac{\beta}{\alpha}\operatorname*{Id}=\left(1-\frac{1}{2}\left(1+\frac{\beta}{\alpha}\right)\right)\operatorname*{Id}+\frac{1}{2}\left(1+\frac{\beta}{\alpha}\right)N.$ (41)
An identical expression holds for $R_{B}^{\beta,\alpha}$ with $\beta/\alpha$ replaced by $\alpha/\beta$. It follows that $R_{A}^{\alpha,\beta}$ and $R_{B}^{\beta,\alpha}$ are $\lambda_{\alpha,\beta}$-conic and $\lambda_{\beta,\alpha}$-conic operators, respectively, where
$\lambda_{\alpha,\beta}:=\frac{1}{2}\left(1+\frac{\beta}{\alpha}\right)\qquad\quad\text{and}\qquad\quad\lambda_{\beta,\alpha}:=\frac{1}{2}\left(1+\frac{\alpha}{\beta}\right),$ (42)
and we conclude that (40) indeed is an averaged iteration of a composition of conic operators. Using that an operator is firmly nonexpansive if and only if it is the resolvent of a maximally monotone operator [BauschkeCombettes2017, Corollary 23.9], we obtain the following characterization of the relaxed resolvents.
Proposition 7.1:
Let $\mathcal{C}_{\lambda}$ be the class of $\lambda$-conic operators $C:\mathcal{H}\to\mathcal{H}$. Then
$\mathcal{C}_{\lambda_{\alpha,\beta}}$ $=\{R:\mathcal{H}\to\mathcal{H}\mid R=R_{A}^{\alpha,\beta}\text{ for some }A\in\mathcal{A}\},$
$\mathcal{C}_{\lambda_{\beta,\alpha}}$ $=\{R:\mathcal{H}\to\mathcal{H}\mid R=R_{B}^{\beta,\alpha}\text{ for some }B\in\mathcal{A}\},$
where $\lambda_{\alpha,\beta}$ and $\lambda_{\beta,\alpha}$ are given in (42), and $R_{A}^{\alpha,\beta}$ and $R_{B}^{\beta,\alpha}$ are defined in (39).
Let us examine the conic constants $\lambda_{\alpha,\beta}$ and $\lambda_{\beta,\alpha}$ more closely. Since $\alpha,\beta>0$, we immediately have $\lambda_{\alpha,\beta},\lambda_{\beta,\alpha}>\tfrac{1}{2}$. The product of the two constants satisfies
@@
$\lambda_{\alpha,\beta}\lambda_{\beta,\alpha}=\frac{1}{4}\left(1+\frac{\beta}{\alpha}\right)\left(1+\frac{\alpha}{\beta}\right)=\frac{1}{4}\left(2+\frac{\beta}{\alpha}+\frac{\alpha}{\beta}\right)\geq 1$
with equality if and only if $\alpha=\beta$. Moreover, $\lambda_{\alpha,\beta}<1$ if and only if $\lambda_{\beta,\alpha}>1$, and vice versa. Together with ˜7.1, this implies that exactly one of the following holds:
  1. i)
    $R_{A}^{\alpha,\beta}$ and $R_{B}^{\beta,\alpha}$ are nonexpansive implying that $R_{B}^{\beta,\alpha}R_{A}^{\alpha,\beta}$ is nonexpansive (if $\alpha=\beta$),
  2. ii)
    $R_{A}^{\alpha,\beta}$ is $\lambda_{\alpha,\beta}$-conic, $R_{B}^{\beta,\alpha}$ is $\lambda_{\beta,\alpha}$-averaged, and there exists $(A,B)\in\mathcal{A}\times\mathcal{A}$ such that $R_{B}^{\beta,\alpha}R_{A}^{\alpha,\beta}$ is not conic (if $\alpha<\beta$),
  3. iii)
    $R_{A}^{\alpha,\beta}$ is $\lambda_{\alpha,\beta}$-averaged, $R_{B}^{\beta,\alpha}$ is $\lambda_{\beta,\alpha}$-conic, and there exists $(A,B)\in\mathcal{A}\times\mathcal{A}$ such that $R_{B}^{\beta,\alpha}R_{A}^{\alpha,\beta}$ is not conic (if $\alpha>\beta$).
When $\alpha=\beta$ and $\theta\in(0,2)$, as in the unconditionally convergent case for maximally monotone $A$ and $B$, the relaxed resolvents are nonexpansive and algorithm (40) reduces to a $\tfrac{\theta}{2}$-averaged fixed-point iteration of the composition of the two nonexpansive operators. Consequently, the algorithm operator $T_{A,B}^{\alpha,\beta,\theta}$ is averaged. It is straightforward to verify that an operator is averaged if and only if it can be written as an averaging of a conic operator. The above discussion therefore lets us paraphrase [ryu2020uniqueness, Theorem 2] as follows.
The algorithm in (38) converges unconditionally over $\mathcal{A}\times\mathcal{A}$ if and only if its associated algorithm operator $T_{A,B}^{\alpha,\beta,\theta}$ in (40) is an averaged operator for all $(A,B)\in\mathcal{A}\times\mathcal{A}$.
In the optimization setting, our analysis shows that the monotone-operator parameter requirement $\alpha=\beta$ can be relaxed to the inequality $\tfrac{\alpha}{\beta}>\tfrac{\theta}{2}$ (see (5)), meaning that $\alpha$ must be sufficiently large relative to $\beta$. Whenever $\alpha\neq\beta$, there exist operators $(A,B)\in\mathcal{A}\times\mathcal{A}$—and even $(A,B)\in\partial\mathcal{F}\times\partial\mathcal{F}$, as we will demonstrate later—for which the algorithm operator $T_{A,B}^{\alpha,\beta,\theta}$ is not averaged although the corresponding algorithm in (38) converges unconditionally over $\partial\mathcal{F}\times\partial\mathcal{F}$. This implies that alternative proof techniques are required—such as the Lyapunov analysis developed in this paper—to establish convergence in the convex optimization setting. See Figure˜3 for two-dimensional graphical illustrations of the potential lack of averagedness and Figure˜4 for similar illustrations overlaid with an example. Illustrations similar to those in Figures˜2, 3 and 4 have appeared before in the literature as a tool for understanding algorithm dynamics [Eckstein1989, GiselssonMoursi2021Compositions, RyuHannahYin2022SRG].
$^{\star}zzzz^{\star}$$zz^{\star}$
Figure 3: Two-dimensional representations of the possible locations of $z^{+}\in\mathbb{R}^{2}$ in (40) relative to an arbitrary fixed point $z^{\star}\in\mathbb{R}^{2}$, for a given starting point $z\in\mathbb{R}^{2}$. Left figure: $(\alpha,\beta,\theta)=(1,9,0.2)$ implying $(\lambda_{\alpha,\beta},\lambda_{\beta,\alpha})=(5,0.56)$, i.e., that $z^{+}$ is obtained by applying to $z$ a $0.18$-averaging of a composition of a 55-conic and a $0.56$-conic operator. Middle figure: $(\alpha,\beta,\theta)=(1,0.2,1.5)$ implying $(\lambda_{\alpha,\beta},\lambda_{\beta,\alpha})=(0.6,3)$, i.e., that $z^{+}$ is obtained by applying to $z$ a $0.25$-averaging of a composition of a $0.6$-conic and a 33-conic operator. Right figure: $(\alpha,\beta,\theta)=(1,1,1.5)$ implying $(\lambda_{\alpha,\beta},\lambda_{\beta,\alpha})=(1,1)$, i.e., that $z^{+}$ is obtained by applying to $z$ a $0.75$-averaging of a composition of two nonexpansive operators. The left and middle figures illustrate that, in these regimes, the associated algorithm operators may fail to be averaged.
$zz^{\star}$$C_{1}$$C_{2}$$zz^{\star}$$r_{1}$$r_{2}$$^{+}zz$$z^{\star}$$C_{1}$$C_{2}$$zz^{\star}$$r_{1}$$r_{2}$$z^{+}$
Figure 4: Two-dimensional illustration of one step of (40) applied to the feasibility problem (43), with the regions from Figure˜3 indicating where $z^{+}$ may lie shown in the background. We denote $r_{1}=R_{\partial\iota_{C_{1}}}^{\alpha,\beta}(z)$, $r_{2}=R_{\partial\iota_{C_{2}}}^{\beta,\alpha}(r_{1})$, and $z^{+}=\left(1-\frac{\theta\beta}{\alpha+\beta}\right)z+\frac{\theta\beta}{\alpha+\beta}r_{2}$. Left figure: $(\alpha,\beta,\theta)=(1,3.5,0.45)$ implying $\frac{\theta\beta}{\alpha+\beta}=0.35$ and $(\lambda_{\alpha,\beta},\lambda_{\beta,\alpha})=(2.25,0.64)$. Thus, the first relaxed resolvent is conic but not nonexpansive, and the second is averaged. The resulting $z^{+}$ lies outside the nonexpansiveness circle. Right figure: $(\alpha,\beta,\theta)=(1,1,1.5)$ implying $\frac{\theta\beta}{\alpha+\beta}=0.75$ and $(\lambda_{\alpha,\beta},\lambda_{\beta,\alpha})=(1,1)$, i.e., the Douglas–Rachford setting. In this case, both relaxed resolvents are nonexpansive reflected resolvents. Consequently, the point $z^{+}$ remains inside the nonexpansiveness circle. Both these algorithms converge, although the distance to a fixed point may temporarily increase between iterations in the left example.
Let us present the problem behind the illustration in Figure˜4. Let $c\in\mathbb{R}\setminus\{0\}$ and consider $f=\iota_{C_{1}}$ and $g=\iota_{C_{2}}$, where
$C_{1}=\{x\in\mathbb{R}^{2}\mid x=(ct,t)\text{ for }t\in\mathbb{R}\}\quad\text{and}\quad C_{2}=\{x\in\mathbb{R}^{2}\mid x=(0,t)\text{ for }t\in\mathbb{R}\}.$
The problem of finding an $x\in\mathbb{R}^{2}$ such that
$0\in\partial\iota_{C_{1}}(x)+\partial\iota_{C_{2}}(x)$ (43)
(where $\partial\iota_{C_{i}}$ is the normal cone of $C_{i}\subset\mathbb{R}^{2}$) is a convex feasibility problem whose unique solution is $x=(0,0)$. Figure˜4 shows that the algorithm operator $T_{A,B}^{\alpha,\beta,\theta}$ can fail to be averaged for this problem by visualizing parameter choices for which $T_{\partial\iota_{C_{1}},\partial\iota_{C_{2}}}^{\alpha,\beta,\theta}$ is convergent yet neither averaged nor nonexpansive. The example uses $c=\tfrac{1}{2}$. In this case, the algorithm operator is represented by the matrix:
$T:=T_{\partial\iota_{C_{1}},\partial\iota_{C_{2}}}^{\alpha,\beta,\theta}=\begin{bmatrix}1-\frac{\theta}{5}&-\frac{2\theta}{5}\\ \frac{2\theta\beta}{5\alpha}&1-\frac{\theta\beta}{5\alpha}\end{bmatrix}$.
Because (43) is a feasibility problem, the resolvent steps reduce to projections, so the algorithm depends on $\alpha$ and $\beta$ only through the ratio $\beta/\alpha$. The iteration matrix $T$ is invertible for every choice of $(\alpha,\beta,\theta)$ satisfying the parameter restrictions in (5), and convergence to the unique fixed point at $(0,0)$ follows from an eigenvalue analysis. The asymptotic rate of convergence is governed by the magnitude of the largest eigenvalue of $T$. Moreover, the largest singular value of $T$ is the (smallest) Lipschitz constant (in $\mathbb{R}^{2}$), which, for problems over $\mathcal{A}\times\mathcal{A}$, is illustrated in Figures˜3 and 4 as the point in the shaded region that lies the farthest from the fixed point.
In the Douglas–Rachford case (with $\theta=1$ corresponding to $\tfrac{1}{2}$-averaging), the magnitude of the largest eigenvalue is known to equal the cosine of the Friedrichs angle [BauschkeEtAl2014FriedrichsAngle], which here is $2/\sqrt{5}\approx 0.89$. This is the best possible rate in the Douglas–Rachford setting with $\theta\in(0,2)$ and $\alpha=\beta$. Since the matrix $T$ is normal in this case, the largest singular value coincides with the magnitude of the largest eigenvalue. Indeed, the right-most illustrations in Figures˜3 and 4 indicate that this is not a coincidence as the set of possible algorithm operators must have Lipschitz constant at most 1.
We can optimize the magnitude of the largest eigenvalue of $T$ with two feasible parameter choices
$(\alpha,\beta,\theta)=\left(1,\frac{1-\epsilon}{\sigma},2\sigma\right)\qquad\quad\text{and}\quad\qquad(\alpha,\beta,\theta)=\left(1,\frac{\sigma}{1-\epsilon},2(1-\epsilon)\right),$
where $\sigma=9-4\sqrt{5}\approx 0.056$ and $\epsilon\in(0,1)$. We note that the resulting matrices are unitarily similar with unitary matrix $P\in\mathbb{R}^{2\times 2}$ that has zero on the diagonal and $P_{12}=1$ and $P_{21}=-1$. They therefore share the same eigenvalues and singular values. As $\epsilon\downarrow 0$, we approach the boundary of the feasible parameter region. For small enough $\epsilon>0$, the largest eigenvalue magnitude of $T$ becomes arbitrarily close to $\sqrt{(3+2\sigma)/5}\approx 0.79$ in both cases, which is a strictly better rate than in the Douglas–Rachford case. The largest singular values become in both cases arbitrarily close to $\sqrt{1+2(1-\sigma)(\sqrt{\sigma^{2}+4}-\sigma)/5}\approx 1.32$. A largest singular value exceeding 1 is equivalent to the operator being not nonexpansive, and therefore not averaged. This shows that indeed, the algorithm operator can fail to be averaged and still converge unconditionally. In the left illustration of Figure˜4, we showcase one parameter choice for which this is the case.
To characterize the broader class of unconditionally convergent methods in the convex optimization setting, we had to step outside the usual framework of averaged operators. While allowing iterations that are not averaged is not, by itself, the fundamental source of the improved convergence observed in our example, it does suggest that the additional freedom in parameter selection can translate into meaningful performance gains.
@@ 8 Conclusions
In this work, we have shown that, in the convex optimization setting, the Douglas–Rachford splitting method is not unique as an unconditionally convergent, frugal, no-lifting resolvent-splitting scheme. When the operators are subdifferentials of proper, closed, convex functions, we identify a strictly larger class of such methods and provide a complete, sharp characterization of all parameter choices that yield unconditional convergence. This characterization, in turn, immediately yields new, provably convergent families of ADMM- and Chambolle–Pock–type algorithms via their Douglas–Rachford reformulations, including parameter regimes that were previously not known to converge.
@@
Appendix A Verification of Equality (LE) in Proposition˜4.1
The equality in (LE) is also verified symbolically in [nilsson2025verification].
@@
Expanding the norms of the quantities in Definition˜4.2, we obtain the following expressions in terms of quadratic forms for all $k\in\mathbb{N}$:
$V^{(1)}_{k}$ $=\mathcal{Q}\left(\underbrace{\begin{bmatrix}\theta-\tau+1&-\frac{\theta(1+\tau)}{2}&0&0&\tau-1-\frac{\theta}{2}\\[6.0pt] *&\theta\tau&0&0&\frac{\theta}{2}\\[6.0pt] *&*&0&0&0\\[6.0pt] *&*&*&0&0\\[6.0pt] *&*&*&*&1\end{bmatrix}}_{=:Q_{V}^{(1)}},\begin{bmatrix}x^{k}_{1}-x^{\star}\\[6.0pt] x^{k}_{2}-x^{\star}\\[6.0pt] x^{k+1}_{1}-x^{\star}\\[6.0pt] x^{k+1}_{2}-x^{\star}\\[6.0pt] z^{k}-z^{\star}\end{bmatrix}\right)+2\alpha(1-\tau)D_{f}(x^{k}_{1},x^{\star},u^{\star}),$
$V^{(2)}_{k}$ $=\mathcal{Q}\left(\underbrace{\begin{bmatrix}\theta+\tau-1&-\frac{\theta(1+\tau)}{2}&0&0&-\frac{\theta}{2}\\[6.0pt] *&\theta\tau&0&0&\frac{\theta}{2}\\[6.0pt] *&*&0&0&0\\[6.0pt] *&*&*&0&0\\[6.0pt] *&*&*&*&1\end{bmatrix}}_{=:Q_{V}^{(2)}},\begin{bmatrix}x^{k}_{1}-x^{\star}\\[6.0pt] x^{k}_{2}-x^{\star}\\[6.0pt] x^{k+1}_{1}-x^{\star}\\[6.0pt] x^{k+1}_{2}-x^{\star}\\[6.0pt] z^{k}-z^{\star}\end{bmatrix}\right)+2\alpha(\tau-1)D_{f}(x^{\star},x^{k}_{1},u^{k}_{1}),$
@@
$\qquad+2\alpha(\tau-1)D_{f}(x_{1}^{k+1},x_{1}^{k},u_{1}^{k}),$
and
$V_{k+1}^{(1)}$ $=\mathcal{Q}\left(Q_{V}^{(1)},\begin{bmatrix}x^{k+1}_{1}-x^{\star}\\ x^{k+1}_{2}-x^{\star}\\ x^{k+2}_{1}-x^{\star}\\ x^{k+2}_{2}-x^{\star}\\ z^{k+1}-z^{\star}\end{bmatrix}\right)+2\alpha(1-\tau)D_{f}(x^{k+1}_{1},x^{\star},u^{\star})$
$V_{k+1}^{(2)}$ $=\mathcal{Q}\left(Q_{V}^{(2)},\begin{bmatrix}x^{k+1}_{1}-x^{\star}\\ x^{k+1}_{2}-x^{\star}\\ x^{k+2}_{1}-x^{\star}\\ x^{k+2}_{2}-x^{\star}\\ z^{k+1}-z^{\star}\end{bmatrix}\right)+2\alpha(\tau-1)D_{f}(x^{\star},x^{k+1}_{1},u^{k+1}_{1}).$
For a matrix $Q\in\mathbb{S}^{5}$ with structure $Q=\begin{bmatrix}q_{11}&q_{12}&0&0&q_{13}\\ *&q_{22}&0&0&q_{23}\\ *&*&0&0&0\\ *&*&*&0&0\\ *&*&*&*&1\end{bmatrix}$ we have that
@@
where the first equality is due to the zero columns of $Q$ and the second equality follows from the update step of Algorithm˜1. Therefore
$V_{k+1}^{(1)}$ $=\mathcal{Q}\left(\underbrace{\begin{bmatrix}\theta^{2}&-\theta^{2}&-\theta(\tau-1-\frac{\theta}{2})&-\frac{\theta^{2}}{2}&-\theta\\[6.0pt] *&\theta^{2}&\theta(\tau-1-\frac{\theta}{2})&\frac{\theta^{2}}{2}&\theta\\[6.0pt] *&*&1-\tau+\theta&-\frac{\theta(1+\tau)}{2}&\tau-1-\frac{\theta}{2}\\[6.0pt] *&*&*&\theta\tau&\frac{\theta}{2}\\[6.0pt] *&*&*&*&1\end{bmatrix}}_{=M^{\top}Q_{V}^{(1)}M},\begin{bmatrix}x^{k}_{1}-x^{\star}\\[6.0pt] x^{k}_{2}-x^{\star}\\[6.0pt] x^{k+1}_{1}-x^{\star}\\[6.0pt] x^{k+1}_{2}-x^{\star}\\[6.0pt] z^{k}-z^{\star}\end{bmatrix}\right)+2\alpha(1-\tau)D_{f}(x^{k+1}_{1},x^{\star},u^{\star})$
and
$V_{k+1}^{(2)}$ $=\mathcal{Q}\left(\underbrace{\begin{bmatrix}\theta^{2}&-\theta^{2}&\frac{\theta^{2}}{2}&-\frac{\theta^{2}}{2}&-\theta\\[6.0pt] *&\theta^{2}&-\frac{\theta^{2}}{2}&\frac{\theta^{2}}{2}&\theta\\[6.0pt] *&*&\theta+\tau-1&-\frac{\theta(1+\tau)}{2}&-\frac{\theta}{2}\\[6.0pt] *&*&*&\theta\tau&\frac{\theta}{2}\\[6.0pt] *&*&*&*&1\end{bmatrix}}_{=M^{\top}Q_{V}^{(2)}M},\begin{bmatrix}x^{k}_{1}-x^{\star}\\[6.0pt] x^{k}_{2}-x^{\star}\\[6.0pt] x^{k+1}_{1}-x^{\star}\\[6.0pt] x^{k+1}_{2}-x^{\star}\\[6.0pt] z^{k}-z^{\star}\end{bmatrix}\right)+2\alpha(\tau-1)D_{f}(x^{\star},x^{k+1}_{1},u^{k+1}_{1}).$
By Lemma˜2.1 we get that
@@
and
$D_{f}(x^{\star},x^{k}_{1},u^{k}_{1})$ $-D_{f}(x^{\star},x^{k+1}_{1},u^{k+1}_{1})-D_{f}(x^{k+1}_{1},x^{k}_{1},u^{k}_{1})=\langle u^{k+1}_{1}-u^{k}_{1},x^{\star}-x^{k+1}_{1}\rangle$
$=-\frac{1}{\alpha}\langle z^{k+1}-z^{k}-(x^{k+1}_{1}-x^{k}_{1}),x^{k+1}_{1}-x^{\star}\rangle$
$=-\frac{1}{\alpha}\langle\theta(x^{k}_{2}-x^{k}_{1})-(x^{k+1}_{1}-x^{k}_{1}),x^{k+1}_{1}-x^{\star}\rangle$
[…]
Using these equalities we find that
@@
and similarly
[…]
$\quad\;+2\alpha(\tau-1)\left(D_{f}(x^{\star},x^{k}_{1},u^{k}_{1})-D_{f}(x^{\star},x^{k+1}_{1},u^{k+1}_{1})-D_{f}(x^{k+1}_{1},x^{k}_{1},u^{k}_{1})\right)$
[…]
This gives quadratic expressions for the left hand side of equation ˜LE. We also note that the right hand side of ˜LE can be expressed as
$\theta\alpha I_{k}$ $=\theta\alpha\left(\langle u_{1}^{k}-u^{\star},x_{1}^{k}-x^{\star}\rangle+\langle u_{2}^{k}+u^{\star},x_{2}^{k}-x^{\star}\rangle+\langle u_{1}^{k+1}-u^{\star},x_{1}^{k+1}-x^{\star}\rangle+\langle u_{2}^{k+1}+u^{\star},x_{2}^{k+1}-x^{\star}\rangle\right)$
@@
$Q_{V}^{(i)}-M^{\top}Q_{V}^{(i)}M-Q_{R}^{(i)}+2(1-\tau)Q^{(i)}$ $=\theta Q_{I},$
where $Q_{I}$ is given by (44), for $i\in\{1,2\}$:
$Q_{V}^{(1)}-M^{\top}Q_{V}^{(1)}M-Q_{R}^{(1)}+2(1-\tau)Q^{(1)}$ $=\begin{bmatrix}\theta-\tau+1&-\frac{\theta(1+\tau)}{2}&0&0&\tau-1-\frac{\theta}{2}\\[6.0pt] *&\theta\tau&0&0&\frac{\theta}{2}\\[6.0pt] *&*&0&0&0\\[6.0pt] *&*&*&0&0\\[6.0pt] *&*&*&*&1\end{bmatrix}$

### v2

@@ 6.4 Extending Generalized Alternating Projections
$R_{\alpha c\operatorname{dist}_{C}^{2}}^{\alpha,\beta}$ $=\left(1+\frac{\beta}{\alpha}\right){\rm{prox}}_{\alpha c\operatorname{dist}_{C}^{2}}-\frac{\beta}{\alpha}\operatorname*{Id}$
$=\left(1+\frac{\beta}{\alpha}\right)\left(\frac{c^{-1}}{c^{-1}+2\alpha}\operatorname*{Id}+\frac{2\alpha}{c^{-1}+2\alpha}\Pi_{C}\right)-\frac{\beta}{\alpha}\operatorname*{Id}$
$=\left(1-\frac{2(\alpha+\beta)}{c^{-1}+2\alpha}\right)\operatorname*{Id}+\frac{2(\alpha+\beta)}{c^{-1}+2\alpha}\Pi_{C}.$
The expression $\frac{2(\alpha+\beta)}{c^{-1}+2\alpha}$ is increasing for $c>0$ and $\frac{2(\alpha+\beta)}{c^{-1}+2\alpha}\in\left(0,1+\frac{\beta}{\alpha}\right]$ for $c\in(0,+\infty]$. Let us introduce $\mu_{1}\coloneqq\frac{2(\alpha+\beta)}{c^{-1}+2\alpha}\in\left(0,1+\frac{\beta}{\alpha}\right]$ and $\mu_{2}\coloneqq\frac{2(\alpha+\beta)}{d^{-1}+2\beta}\in\left(0,1+\frac{\alpha}{\beta}\right]$, where $\mu_{1}$ and $\mu_{2}$ can take any values in these intervals for appropriate choices of $c$ and $d$ respectively, and define
$R_{C}^{\mu_{1}}$ $=(1-\mu_{1})\operatorname*{Id}+\mu_{1}\Pi_{C}\qquad\text{and}\qquad R_{D}^{\mu_{2}}=(1-\mu_{2})\operatorname*{Id}+\mu_{2}\Pi_{D}.$
Applying our Douglas–Rachford algorithm reformulation in (7) to this problem formulation gives the iteration
$z^{k+1}$ $=\left(1-\frac{\theta\beta}{\alpha+\beta}\right)z^{k}+\frac{\theta\beta}{\alpha+\beta}R_{D}^{\mu_{2}}R_{C}^{\mu_{1}}z^{k}=\left(1-\lambda\right)z^{k}+\lambda R_{D}^{\mu_{2}}R_{C}^{\mu_{1}}z^{k},$
@@
where $\lambda:=\frac{\theta\beta}{\alpha+\beta}$. Now, by Theorem˜3.3 we have convergence if $\theta\in\left(0,\min\{2,2\alpha/\beta\}\right)$, implying that
$\lambda\in\left(0,\min\left\{\frac{2\beta}{\alpha+\beta},\frac{2\alpha}{\alpha+\beta}\right\}\right).$
Let us also introduce $\gamma\coloneqq\beta/\alpha>0$, to get that Algorithm˜6 converges if
$\mu_{1}\in(0,1+\gamma]\qquad\text{and}\qquad\mu_{2}\in(0,1+\gamma^{-1}]\qquad\text{and}\qquad\lambda\in\left(0,\min\left\{\frac{2}{1+\gamma},\frac{2}{1+\gamma^{-1}}\right\}\right).$
@@ 7 Conclusions
In this work, we have shown that, in the convex optimization setting, the Douglas–Rachford splitting method is not the unique unconditionally convergent, frugal, no-lifting resolvent-splitting scheme. When the operators are subdifferentials of proper, closed, convex functions, we identify a strictly larger class of such methods and provide a complete, sharp characterization of all parameter choices that yield unconditional convergence. This characterization, in turn, immediately yields new, provably convergent families of ADMM- and Chambolle–Pock–type algorithms via their Douglas–Rachford reformulations, including parameter regimes that were previously not known to converge.
@@
Appendix A Verification of Equality (VR) in Lemma˜4.1 and (LE) in Proposition˜4.1
@@
A.1 Verification of (VR) in Lemma˜4.1
Let $\bar{V}_{k}$ denote the right hand side of the first equality of (VR). Expanding the norms of the quantities in Definition˜4.2 and Lemma˜4.1 gives the following expressions in terms of quadratic forms for all $k\in\mathbb{N}$,
$V_{k}$ $=\mathcal{Q}\left(\underbrace{\begin{bmatrix}\theta-\tau+1&-\frac{\theta(1+\tau)}{2}&0&0&\tau-1-\frac{\theta}{2}\\[6.0pt] *&\theta\tau&0&0&\frac{\theta}{2}\\[6.0pt] *&*&0&0&0\\[6.0pt] *&*&*&0&0\\[6.0pt] *&*&*&*&1\end{bmatrix}}_{=:Q_{V}^{(1)}},\begin{bmatrix}x^{k}_{1}-x^{\star}\\[6.0pt] x^{k}_{2}-x^{\star}\\[6.0pt] x^{k+1}_{1}-x^{\star}\\[6.0pt] x^{k+1}_{2}-x^{\star}\\[6.0pt] z^{k}-z^{\star}\end{bmatrix}\right)+2\alpha(1-\tau)D_{f}(x^{k}_{1},x^{\star},u^{\star}),$
$\bar{V}_{k}$ $=\mathcal{Q}\left(\underbrace{\begin{bmatrix}\theta+\tau-1&-\frac{\theta(1+\tau)}{2}&0&0&-\frac{\theta}{2}\\[6.0pt] *&\theta\tau&0&0&\frac{\theta}{2}\\[6.0pt] *&*&0&0&0\\[6.0pt] *&*&*&0&0\\[6.0pt] *&*&*&*&1\end{bmatrix}}_{=:Q_{V}^{(2)}},\begin{bmatrix}x^{k}_{1}-x^{\star}\\[6.0pt] x^{k}_{2}-x^{\star}\\[6.0pt] x^{k+1}_{1}-x^{\star}\\[6.0pt] x^{k+1}_{2}-x^{\star}\\[6.0pt] z^{k}-z^{\star}\end{bmatrix}\right)+2\alpha(\tau-1)D_{f}(x^{\star},x^{k}_{1},u^{k}_{1})$
$=\mathcal{Q}\left(Q_{V}^{(1)},\begin{bmatrix}x^{k}_{1}-x^{\star}\\[6.0pt] x^{k}_{2}-x^{\star}\\[6.0pt] x^{k+1}_{1}-x^{\star}\\[6.0pt] x^{k+1}_{2}-x^{\star}\\[6.0pt] z^{k}-z^{\star}\end{bmatrix}\right)-2(\tau-1)\langle x_{1}^{k}-x^{\star},\underbrace{z^{k}-z^{\star}-(x_{1}^{k}-x^{\star})}_{=\alpha(u_{1}^{k}-u^{\star})}\rangle+2\alpha(\tau-1)D_{f}(x^{\star},x^{k}_{1},u^{k}_{1}).$
By Definition˜4.1 and by Lemma˜2.2 we then have that
$\bar{V}_{k}-V_{k}=-2\alpha(\tau-1)\langle x_{1}^{k}-x^{\star},u_{1}^{k}-u^{\star}\rangle+2\alpha(\tau-1)\left(D_{f}(x_{1}^{k},x^{\star},u^{\star})+D_{f}(x^{\star},x_{1}^{k},u_{1}^{k})\right)=0.$
Similarly, if we let $\bar{R}_{k}$ denote the right hand side of the second equality of (VR), we obtain the following expressions in terms of quadratic forms for all $k\in\mathbb{N}$,
@@
$\qquad+2\alpha(\tau-1)D_{f}(x_{1}^{k+1},x_{1}^{k},u_{1}^{k}).$
Moreover
$Q_{R}^{(2)}-Q_{R}^{(1)}$ $=(\tau-1)\begin{bmatrix}2-2\theta&\theta&\theta-2&0&0\\ *&0&-\theta&0&0\\ *&*&2&0&0\\ *&*&*&0&0\\ *&*&*&*&0\end{bmatrix}$. (40)
By Definition˜4.1 and by Lemma˜2.2 we have that the difference of the Bregman terms of $\bar{R}_{k}$ and $R_{k}$ is
$2\alpha(\tau-1)\left(D_{f}(x_{1}^{k+1},x_{1}^{k},u_{1}^{k})+D_{f}(x_{1}^{k},x_{1}^{k+1},u_{1}^{k+1})\right)=\langle x_{1}^{k}-x_{1}^{k+1},u_{1}^{k}-u_{1}^{k+1}\rangle$
$=2(\tau-1)\langle x_{1}^{k}-x_{1}^{k+1},z^{k}-x_{1}^{k}-(z^{k}+\theta(x_{2}^{k}-x_{1}^{k})-x_{1}^{k+1})\rangle$
$=2(1-\tau)\langle x_{1}^{k}-x_{1}^{k+1},x_{1}^{k}-x_{1}^{k+1}+\theta(x_{2}^{k}-x_{1}^{k})\rangle$
[…]
Together with (40) this shows that $\bar{R}_{k}-R_{k}=0$.
A.2 Verification of Equality (LE) in Proposition˜4.1
What remains to show is the Lyapunov equality ˜LE. We use the quadratic expressions from Section˜A.1. We have for all $k\in\mathbb{N}$ that
$V_{k+1}$ $=\mathcal{Q}\left(Q_{V}^{(1)},\begin{bmatrix}x^{k+1}_{1}-x^{\star}\\ x^{k+1}_{2}-x^{\star}\\ x^{k+2}_{1}-x^{\star}\\ x^{k+2}_{2}-x^{\star}\\ z^{k+1}-z^{\star}\end{bmatrix}\right)+2\alpha(1-\tau)D_{f}(x^{k+1}_{1},x^{\star},u^{\star}).$
Since the third and fourth rows and columns of $Q_{V}^{(1)}$ are zero, we get that the quadratic form in $V_{k+1}$ can be expressed as follows:
@@
where the first equality is due to the zero columns of $Q_{V}^{(1)}$ and the second equality follows from the update step of Algorithm˜1.
For the Bregman-type distances we have by Lemma˜2.1 that
@@
Combining this equality with our expressions for $V_{k}$, $V_{k+1}$ and $R_{k}$ we get for each $k\in\mathbb{N}$ that
@@
This gives a quadratic expression for $V_{k}-V_{k+1}-R_{k}$. Note also that the remaining term of ˜LE, $\theta\alpha I_{k}$, can be expressed as
$\theta\alpha I_{k}$ $=$
$\theta\alpha\left(\langle u_{1}^{k}-u^{\star},x_{1}^{k}-x^{\star}\rangle+\langle u_{2}^{k}+u^{\star},x_{2}^{k}-x^{\star}\rangle+\langle u_{1}^{k+1}-u^{\star},x_{1}^{k+1}-x^{\star}\rangle+\langle u_{2}^{k+1}+u^{\star},x_{2}^{k+1}-x^{\star}\rangle\right)$
@@
$Q_{V}^{(1)}-M^{\top}Q_{V}^{(1)}M-Q_{R}^{(1)}+2(1-\tau)Q_{D}$ $=\theta Q_{I},$
where $Q_{I}$ is given by (41):
$Q_{V}^{(1)}-M^{\top}Q_{V}^{(1)}M-Q_{R}^{(1)}+2(1-\tau)Q_{D}$ $=\begin{bmatrix}\theta-\tau+1&-\frac{\theta(1+\tau)}{2}&0&0&\tau-1-\frac{\theta}{2}\\[6.0pt] *&\theta\tau&0&0&\frac{\theta}{2}\\[6.0pt] *&*&0&0&0\\[6.0pt] *&*&*&0&0\\[6.0pt] *&*&*&*&1\end{bmatrix}$